OpenStudy (anonymous):

write the expression as a complex in standard form (1+i)+(5+2i)=

OpenStudy (sonic):

6+3i

OpenStudy (sonic):

3i+6 I mean. -.-

OpenStudy (amistre64):

I think its real + imaginary. 6 + 3i was correct :)

OpenStudy (amistre64):

i is treated like any other variable. the only exception is that i^2 = -1.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

write the expression as a complex number in standard form (1+i)(-1+i)

OpenStudy (amistre64):

1(-1) +1(i) -1(i) + i(i) -1 + i^2 -1 +(-1) = -2 -2 + 0i or simply exclude the "imaginary" part: -2